OpenStudy (anonymous):

Solve 3 over x minus 3 equals the quantity of x over x minus 3, minus three-halves for x and determine if the solution is extraneous or not. x = 3, non-extraneous x = 3, extraneous x = -3, extraneous x = -3, non-extraneous

OpenStudy (anonymous):

\[\frac{ 3 }{ x-3}=\frac{ x }{ x-3 }-\frac{ 3 }{ 2 }x \] \[\frac{ 3 }{ x-3 }=\frac{ 2x-3x(x-3) }{ 2\left( x-3 \right) }\] multiply both sides by 2(x-3) \[6=2x-3x^2+9x,3x^2-11x+6=0\] 3x^2-9x-2x+6=0 3x(x-3)-2(x-3)=0 (x-3)(3x-2)=0 \[x-3=0 ~or~ x=3~is~an~extraneous~ root.\] or 3x-2=0 calculate x.